OpenStudy (anonymous):

Jaun jogged around the mall 4 times. How far did he jog if the distance of the inside of the mall is 12/16 mi.

OpenStudy (anonymous):

Assuming the mall is circular, the radius of the mall is 12/16 mile or 3/4 mile. The distance around the mall or its circumference is C = 2(pi)(radius) or C = 2(pi)(3/4) = 1.5pi So 4 times around the mall is 4 circumferences, or 4(1.5pi) = 6pi miles. Just from the wording of your question, I assumed that the mall was circular.